### v1

# A Survey of AI Methods for Geometry Preparation and Mesh Generation in Engineering Simulation

Abstract
Artificial intelligence is beginning to ease long-standing bottlenecks in the CAD-to-mesh pipeline. This survey reviews recent advances where machine learning aids part classification, mesh quality prediction, and defeaturing. We explore methods that improve unstructured and block-structured meshing, support volumetric parameterizations, and accelerate parallel mesh generation. We also examine emerging tools for scripting automation, including reinforcement learning and large language models. Across these efforts, AI acts as an assistive technology, extending the capabilities of traditional geometry and meshing tools. The survey highlights representative methods, practical deployments, and key research challenges that will shape the next generation of data-driven meshing workflows.
1 Introduction
High-fidelity simulation is now a cornerstone of engineering and scientific workflows. Yet, getting from a 3D model to a simulation-ready mesh remains one of the most time-consuming and frustrating steps in the process. Engineers still spend countless hours identifying parts, simplifying geometry, setting up meshing parameters, and checking mesh quality. These steps are often repeated across similar models. Despite decades of progress in geometry kernels and meshing algorithms, this part of the workflow remains surprisingly manual and error-prone.
At the 2024 International Meshing Roundtable, a panel of researchers and practitioners, including several authors of this paper, discussed whether artificial intelligence could help. The answer was not that AI will replace traditional methods. Rather, AI might help amplify them by learning from past decisions, catching problems early, and automating tedious steps. In this view, AI is a powerful assistant that complements expert knowledge instead of replacing it.
Artificial intelligence (AI) and machine learning (ML) are broad terms that describe systems that learn from data. In recent years, these tools have made significant progress in areas such as image recognition, language understanding, and robotics. They are now beginning to show promise in simulation workflows. Applications include identifying parts in CAD assemblies, predicting where meshes might fail, guiding how geometry should be simplified, and even generating scripts to automate routine tasks.
To help readers navigate this growing space, we present a taxonomy in Table 1. It summarizes the main types of AI and machine learning methods being used in model preparation and meshing. Each category includes a brief description, examples of how it is applied, and representative citations. For readers who are new to this area, the table provides both a quick reference and a roadmap for the rest of the paper.
Table 1: Taxonomy of AI/ML methods applied to CAD model preparation and meshing workflows, showing descriptions, referenced sections, and representative citations.
AI/ML Category Description Referenced Sections Citations
Classical Machine Learning Rule-based statistical learning (SVMs, decision trees, ensembles, surrogates) 3D Part Classification; Mesh Quality Prediction; Defeaturing; Parallel Mesh Generation [IpRegli2006, Jayanti2006, Breiman2001, OwenIMR2019, OwenOSTI2022, OwenIMR2023, Parry2025, ParryIMR2023, sklearn, gamatie2019, ipek2005, garner2025numa, garner2025dist, tsolakis2024]
Deep Learning – Geometric Models Neural networks specialized for geometry (CNNs, GNNs, Transformers, hybrids) 3D Part Segmentation; Mesh Quality Prediction (baseline); Volumetric Parameterizations; Scripting Automation (graph encodings) [lambourne2021brepnet, dupont2022cadops, cha2023machining, lee2023brepgat, wu2024aagnet, zheng2025sfrgnn, park2025dfgat, zhang2024brepmfr, dai2025brepformer, jayaraman2021uv, wang2023hybrid, uwimana2025segmentation, yang2025hybrid, DeepShapeCode, zhang2024brep2seq, wu2020comprehensive]
Generative & Representation Learning Generative shape representations (implicit SDFs, parametric splines, hybrids, diffusion models) Unstructured Meshing (geometry construction); Volumetric Parameterizations (DL-Polycube, DDPM-Polycube, neural integral mapping) [park2019deepsdf, mescheder2019occupancy, peng2020convolutional, mildenhall2021nerf, gropp2020implicit, larios2021deep, novello2022exploring, dong2024neurcadrecon, dong2025neurcross, sharma2020parsenet, williamson2025neural, maruani2024ponq, he2025sparseflex, liao2018deep, chen2022neural, sundararaman2024self, yu2025dlpolycube, yu2025DDPM, Zhan2025]
Reinforce-ment Learning (RL) Agents learn multi-step geometry/meshing actions via trial-and-error Defeaturing (future RL); Unstructured Meshing (quad, Delaunay); Block Structured Meshing (2D, 3D); Volumetric Parameterizations (sequence optimization); Scripting Automation (block decomposition, thin volume reduction, CAD scripting); Language-Driven Scripting (integration with RL) [pan2023reinforcement, TONG2023102109, Thacher2025, diprete2024reinforcement, haarnoja-2018-sac, Patel2021AutoHex, Zhang2025CADRLHex, owen2024cad, zhang2025reinforcement]
Large Language Models (LLMs) Text-to-code synthesis, multimodal assistants, agentic AI (LLM+RL) Language-Driven Scripting; Scripting Automation (integration with GNN/RL) [zhao2023survey, R1, R4, R13, N1, R5, R6, R14, R18, R26, N2, zhang2024brep2seq, zhang2024brepmfr, wu2020comprehensive, fang2025selfevolving]
The range of applications is broad. Classical machine learning methods support tasks such as part classification and mesh quality prediction. Deep learning models designed for geometric data are being used for segmentation, defeaturing, and volumetric parameterization. Reinforcement learning is being tested for guiding multi-step decisions in meshing workflows. Large language models are beginning to assist with scripting, automation, and natural language interaction with CAD tools.
Many of these efforts are still in early stages, but the momentum is clear. AI is beginning to reshape not only how meshes are generated, but also how geometry is prepared, validated, and integrated into simulation pipelines. This paper surveys the current landscape, highlights where AI is making a difference, and outlines the challenges that remain. These include the need for curated datasets, standardized benchmarks, better representations for CAD geometry, and intelligent systems that can support analysts across the entire modeling workflow.
2 3D Part Classification
Complex assemblies frequently include many repeated mechanisms such as bolts, screws, springs, and bearings. Analysts often spend substantial time identifying these parts and transforming them to prepare for analysis; for example, bolted connections may require specific geometric simplifications, specialized meshing, and boundary condition assignment. At the scale of hundreds of bolts per assembly, this manual process becomes tedious and error prone, which motivates automated 3D part classification as a front end to idealization and meshing.
Traditional non-AI approaches rely on manual tagging or rule-based feature recognition. These methods can work for narrowly defined classes but tend to be brittle across diverse CAD sources, modeling conventions, tolerances, and geometric noise, and they scale poorly as category counts and edge cases grow [Jayanti2006, IpRegli2006]. In contrast, AI can learn invariants and decision boundaries from data, tolerating variability in geometry and topology while scaling to many classes. The broader literature spans point-cloud, volumetric, image, and graph-based representations [BRepNet]. Recent B-rep graph methods strengthen the direct use of CAD topology for recognition and segmentation [Colligan2022, Lee2023].
Formally, part classification is cast as supervised learning. Given a training set $\mathcal{D}=\{(\mathbf{x}_{i},\mathbf{y}_{i})\}_{i=1}^{n}$, with feature vectors $\mathbf{x}_{i}$ and labels $\mathbf{y}_{i}$, the objective is to learn a function $f$ such that $\hat{\mathbf{y}}=f(\mathbf{x})$. Foundational work includes curvature-based descriptors with SVMs for manufacturing-process discrimination and early shape benchmarks for engineering parts [IpRegli2006, Jayanti2006]. Deep-learning approaches for CAD classification followed [Qin2014], while graph and B-rep networks now operate directly on faces, edges, and coedges [BRepNet, Colligan2022, Lee2023]. In applied workflows, Owen et al. compute features from curves, surfaces, and volumes available via a modeling kernel [OwenIMR2023, ACIS], and train ensembles of decision trees, specifically Random Forests [Breiman2001] implemented in scikit-learn [sklearn]. Labels are curated from large single-part repositories such as GrabCAD [GrabCAD] and ABC [Koch2019], with additional pre-categorized corpora like MCB [MCB2020]. Comparative studies in this context show that tree ensembles provide strong accuracy, fast inference, and interpretable feature importances relative to neural baselines on tabular B-rep feature sets [OwenIMR2023, OwenOSTI2022].
Methods and implementation have matured into deployable tools. The Cubit® Geometry and Meshing Toolkit integrates part classification and reduction features to drive predefined idealization recipes for meshing and boundary condition setup [CubitWebsite, OwenIMR2023, OwenIMR2019]. Altair HyperMesh extends this trend with AI-driven shape recognition (ShapeAI) that automatically identifies and groups recurring parts such as bolts and fasteners, representing one of the first production uses of ML in FEA preprocessing [AltairHyperMesh, AltairShapeAI]. Siemens has likewise incorporated machine learning into Simcenter NX through shape recognition and similarity search, enabling automated grouping and selection of geometrically related parts [SimcenterAIWorkflow, SiemensNXShapeRecognition]. At present, however, only a handful of commercial preprocessors have deployed such AI capabilities, underscoring both their promise and the still-limited extent of adoption across the broader simulation tool ecosystem.
Looking forward, promising directions include expanding class taxonomies beyond standard hardware, improving cross-CAD robustness, leveraging B-rep-native GNNs [BRepNet, Colligan2022, Lee2023], and coupling classification with downstream steps such as defeaturing and mesh-quality prediction to close the loop in data-driven model preparation.
3 3D Part Segmentation
Segmentation of CAD models refers to the task of assigning semantic labels to regions within a part, typically at the face level. While part classification applies a single label to an entire component (e.g., “bolt” or “flange”), segmentation provides finer detail, identifying geometric features such as holes, fillets, bosses, or functional surfaces. This granularity is essential for tasks in simulation, manufacturing, and process planning, where different regions of a part may require distinct treatment in meshing, defeaturing, or boundary condition application.
Traditionally, segmentation has relied on rule-based systems that detect features using geometric heuristics like curvature, convexity, and adjacency. These approaches can succeed in narrow, well-defined domains, but often struggle to generalize across diverse CAD sources and modeling practices. Manual segmentation, while accurate, is labor-intensive and impractical for large-scale workflows.
Recent advances in deep learning have enabled more robust and scalable segmentation by learning directly from the geometry and topology of B-rep models. Neural architectures such as convolutional neural networks (CNNs), graph neural networks (GNNs), and Transformers are increasingly adapted to CAD representations. These methods can recognize feature patterns without requiring handcrafted rules, generalize across datasets, and support efficient inference in production workflows.
A number of techniques have emerged to exploit B-rep structure. BRepNet [lambourne2021brepnet] introduced the first convolutional framework operating on directed coedges, enabling accurate face labeling. Extensions such as CADOps-Net [dupont2022cadops] predict both feature types and their construction sequences, while Cha et al. [cha2023machining] apply segmentation to machining operations and group coherent regions via connected components. For deployment in constrained environments, lightweight variants based on pruning [gkrispanis2025towards] and transfer learning [gkrispanis2024enhancing] have been proposed.
Graph-based approaches further generalize segmentation by representing faces and edges as attributed graphs. BRepGAT [lee2023brepgat] applies attention-based message passing, and AAGNet [wu2024aagnet] integrates geometric features into an adjacency graph with multi-task learning. Other variants such as SFRGNN-DA [zheng2025sfrgnn] and DFGAT [park2025dfgat] target domain adaptation and custom descriptors for design feature recognition. Meanwhile, transformer-based methods like BrepMFR [zhang2024brepmfr] and BRepFormer [dai2025brepformer] capture both local and global patterns across B-rep structures, enabling effective transfer from synthetic to real-world data.
Hybrid models combine the strengths of different representations. UV-Net [jayaraman2021uv] operates on UV parameter domains using a CNN-GNN hierarchy. DeepFeature [wang2023hybrid] merges rule-based reasoning with learned embeddings, and MeshCAD-Net [uwimana2025segmentation] fuses BRepNet and MeshCNN features for multi-modal learning. Other systems [yang2025hybrid] use primitive decomposition alongside GNNs to manage complex, interactive feature sets.
While most systems remain in the research or prototype stage, early implementations are emerging. BRepNet has been open-sourced and integrated into academic toolkits, and hybrid pipelines show promising results on synthetic benchmarks and curated real-world datasets. Broader deployment in commercial CAD or meshing software is still limited, but progressing.
Looking forward, segmentation will play a foundational role in automating CAD model preparation. It enables fine-grained control over mesh resolution, defeaturing, and physics-based modeling decisions. Future work includes combining segmentation with simulation-aware labels, improving generalization across CAD systems, and integrating segmentation tightly with downstream AI-driven tools for classification, meshing, and analysis.
4 Mesh Quality Prediction
Anticipating mesh quality before meshing is crucial for complex CAD models because early identification of regions likely to yield poor elements prevents costly iteration and focuses cleanup where it matters most. Instead of using meshing as a diagnostic that requires repeated cycles of mesh creation, quality inspection, control adjustments, and remeshing, an ML predictor can estimate local quality directly from B-rep features prior to element generation, surfacing hotspots in advance and guiding sizing and preparation choices. Prior work demonstrates this direction for tetrahedral meshing and positions the approach as algorithm-agnostic across meshers and parameter regimes [OwenIMR2019, OwenOSTI2022, OwenIMR2023]. Parry [Parry2025] presents an extensive B-rep-based study for mesh-quality prediction in defeaturing workflows, with emphasis on ensembles of decision trees and uncertainty estimation to support decision-making. In an earlier IMR research note, Parry compared a Random Forest against a CNN for local quality prediction, reporting better performance from the tree ensemble in this setting [ParryIMR2023].
For each B-rep entity — vertex, curve (edge), and surface (face) — a fixed-length feature vector encodes local geometry and topology within a bounded neighborhood. Example features include curvatures, dihedral angles, proximity to neighboring features, valence, and length and area ratios. A training corpus of CAD parts is meshed under multiple target sizes and parameter settings to obtain localized quality labels near each entity, typically within one to two element edge lengths. Common labels include scaled Jacobian and inradius; definitions and implementations of these metrics are well established in the literature and libraries used by meshing tools [Knupp2000, Knupp2003, Verdict2007].
As with part classification, supervised learning is used: given a dataset $\mathcal{D}=\{(\mathbf{x}_{i},\mathbf{y}_{i})\}_{i=1}^{n}$, the goal is to learn a function $f$ such that $\hat{\mathbf{y}}=f(\mathbf{x})$. Ensembles of decision trees (Random Forests) [Breiman2001], implemented in scikit-learn [sklearn], deliver strong accuracy, fast inference, and interpretable feature importances on tabular B-rep features; Parry’s note finds RF outperforming a CNN baseline for local quality prediction [ParryIMR2023], consistent with results reported in applied workflows [OwenIMR2019, OwenOSTI2022, OwenIMR2023]. The mesh quality predictor is implemented in the Cubit® Geometry and Meshing Toolkit [CubitWebsite] as part of the defeaturing workflow, where it flags low-quality regions prior to meshing and informs subsequent corrective actions, reducing iteration and accelerating time to analysis [OwenIMR2019, OwenIMR2023, CubitWebsite]. Future work includes broader coverage of element types and quality metrics beyond tetrahedra, cross-mesher generalization, uncertainty-aware decision policies, and integration with graph- and B-rep-based neural models for joint geometry–topology reasoning.
5 Defeaturing
CAD models intended for manufacturing often contain small features such as holes, fillets, and engravings. These details are important for fabrication, but they rarely influence the physics of interest and can make meshing difficult by forcing refinement and producing poorly shaped elements. Defeaturing, the process of removing or simplifying such features, is therefore a critical step before simulation. The challenge is that analysts must decide which features can be removed safely, balancing geometry, physics, and meshing requirements. This makes defeaturing one of the most time-consuming and expertise-driven parts of analysis preparation.
Traditional approaches attempt to reduce this burden, but each has limitations. Tolerant meshing methods can bypass fine artifacts by relaxing geometric fidelity [ANSYS2025, MeshGemsCADSurf, Hu2018TetWild, Staten2024Morph], yet they risk discarding features important for fidelity. Manual tools such as CADfix [CADfix2025] and SpaceClaim [SpaceClaim2025] require significant analyst effort and deep experience. What is needed is a way to automate cleanup decisions that is both scalable and trustworthy.
Machine learning provides such an opportunity. By using mesh quality or simulation outcomes as supervision, ML models can learn to recognize problematic regions of geometry and evaluate corrective operations. Mesh quality is particularly attractive as a label because it is fast to compute, general across meshing tools, and provides a practical proxy for simulation performance. This allows thousands of training examples to be generated automatically, creating the data foundation needed for reliable prediction.
Two recent studies illustrate different directions for ML-driven defeaturing. Shinde et al. [Shinde2024] supervised directly on simulation outcomes for sheet metal parts, labeling operations by their effect on quantities of interest in finite element analysis. This produces labels aligned with engineering objectives, but it requires many expensive simulation runs and limits the approach to specific physics. In contrast, Owen et al. [OwenIMR2019] proposed a more general framework driven by mesh quality prediction. Their method first estimates baseline quality for each B-rep entity, as described in Sec. 4. Entities predicted to produce poor elements are flagged, and candidate defeaturing operations in Cubit® such as remove surface, composite surfaces, blunt tangency or collapse curve are then evaluated by separate ML models that predict the post-operation quality. The system ranks both entities and operations, presenting analysts with prioritized cleanup options.
Owen’s implementation [OwenIMR2019] relied on twelve feature models: nine corresponding to specific defeaturing operations and three for unmodified entities. Each model takes as input a fixed-length “expert feature” vector encoding local geometry and topology such as curvature, angles, adjacency, and area and length ratios. Training labels are mesh quality metrics including scaled Jacobian, scaled in-radius, and deviation. These labels are created by executing the CAD operation, meshing the modified geometry, and measuring quality near the affected entities. This process provides explicit evidence of how each operation impacts meshing so that at prediction time the model can recommend the best operation for a given local configuration. Training used the same ensemble decision tree method introduced in Sec. 4, achieving low prediction error across the range of operations. The models were deployed both interactively and in automated greedy workflows, giving novice users a fast starting point while allowing experts to accept, reject, or refine suggested fixes.
This capability has been released in Cubit® and is already in use, but it remains an active R&D effort. Current work is expanding the feature models to cover more operations, improving generalization, and incorporating reinforcement learning. Whereas the current supervised approach predicts static outcomes for single operations, reinforcement learning would enable multi-step procedures that carry out full defeaturing end-to-end. Another direction is incorporating analyst preferences by mining journal files that capture real sequences of cleanup commands. These data could be used to adapt operation rankings to physics context. For example, structural dynamics analysts may favor different cleanup operations than those working in structural mechanics. Together, these efforts point toward defeaturing that is not only automated but also adaptive and informed by both geometry and human expertise.
6 Unstructured Mesh Generation Methods
Defeaturing prepares CAD models for analysis by removing unnecessary complexity, but the next critical step is mesh generation itself. Even on simplified geometry, the robustness and quality of the mesh directly affect the accuracy, efficiency, and convergence of finite element analysis (FEA). Much of the prior discussion has assumed a CAD foundation, where geometry is defined by parametric solids and surfaces constructed in tools such as SolidWorks or Creo. However, many important applications do not begin with CAD at all. Medical imaging, reverse engineering, topology optimization, and scanned parts all provide geometry in implicit, voxel, or point-sampled form rather than as explicit CAD models. In such settings, the domain itself must first be constructed before meshing. This distinction motivates two categories of AI-enhanced approaches: those that construct the geometric domain from raw data, and those that generate unstructured meshes on an existing domain, whether CAD-based or reconstructed.
6.1 Geometry Construction
When CAD is absent, geometry must be built from discrete or implicit data before meshing is possible. Classical methods include Marching Cubes [lorensen1998marching], Dual Contouring, and level-set extraction, which generate watertight triangulated surfaces from voxelized data or distance fields. AI-driven variants improve robustness, feature preservation, and differentiability. DeepSDF and related implicit networks [park2019deepsdf, mescheder2019occupancy, peng2020convolutional, mildenhall2021nerf] represent shapes as continuous neural functions, from which smooth surfaces, normals, and curvature can be derived analytically [gropp2020implicit, larios2021deep, novello2022exploring]. Such neural implicits provide resolution independence and consistent geometry across scales, properties that can directly benefit downstream meshers by reducing boundary noise and preserving sharp features [dong2024neurcadrecon, dong2025neurcross].
Alternative formulations use neural parameterizations, predicting explicit spline or subdivision patches with direct CAD compatibility [sharma2020parsenet, williamson2025neural]. Compared to implicits, these models ease editing and boundary condition specification at the expense of handling complex topology. Hybrid approaches such as PoNQ [maruani2024ponq] and SparseFlex [he2025sparseflex] combine the flexibility of neural fields with the editability of meshes, enabling both high-resolution reconstruction and downstream editing. Differentiable variants of surface extraction, such as Deep Marching Cubes [liao2018deep] or learned Dual Contouring [chen2022neural, sundararaman2024self], relax discrete decisions into smooth surrogates, allowing gradients to flow through geometry reconstruction and making it possible to couple surface generation with optimization or learning objectives.
Although these methods originate largely in computer graphics, their relevance to simulation lies in providing well-conditioned geometric domains when CAD is not available. Applications include meshing of medical imaging data [zhang2016geometric, zhang2013challenges], topology-optimized structures, scanned parts, or any setting where geometry is defined implicitly. By producing smooth, consistent, and differentiable representations of geometry, they improve the reliability of subsequent unstructured mesh generation for FEA.
6.2 Meshing
Once a domain is available—either from CAD or reconstructed using the methods above—the challenge shifts to unstructured mesh generation itself. Classical algorithms such as Delaunay triangulation, advancing front, and paving remain central for tetrahedral, triangular, and quadrilateral meshing. These algorithms are mature and widely deployed, but still rely on heuristic rules for point insertion, front propagation, smoothing, and refinement. Such heuristics can be brittle across diverse geometries and often require manual intervention to achieve high-quality meshes.
AI offers the ability to augment these algorithms by learning adaptive policies that replace or refine heuristics. Reinforcement learning has been applied directly to advancing-front quad meshing: Pan et al. [pan2023reinforcement] formulate front updates as a sequential decision process and train a Soft Actor-Critic (SAC) policy that produces higher-quality meshes than rule-based strategies. Tong et al. [TONG2023102109] extend this with supervised pretraining and RL fine-tuning, improving singularity placement and cross-field alignment—two key challenges in quad meshing. More recently, Thacher, Persson, and Pan [Thacher2025] applied reinforcement learning with graph neural networks to generate then improve connectivity in triangular Delaunay mesh generation. Their method learns point placement and connectivity decisions that yield meshes of comparable or higher quality than established codes such as Triangle or DistMesh, while retaining the theoretical guarantees of Delaunay triangulation. This work demonstrates one of the clearest examples of AI directly enhancing a classical meshing algorithm used in FEA.
Complementary ideas from graphics illustrate how meshing kernels can be embedded into learning frameworks. Neural subdivision methods [liu2020neural, fruhauf2024quadify] retain classical combinatorics while learning residual vertex updates, improving smoothness and feature preservation. Mesh simplification guided by task objectives, such as MeshCNN [hanocka2019meshcnn], shows how decimation can be coupled to downstream goals, a principle analogous to adaptive refinement or coarsening in FEA. Hybrid pipelines, such as Quadify [fruhauf2024quadify], combine classical guarantees with neural refinement, providing a model for CAE workflows where correctness must be preserved but adaptivity is essential.
At present, AI-enhanced unstructured meshing remains at a low technology readiness level. Reinforcement learning for advancing front and Delaunay meshing are the most direct examples with FEA relevance, but they are still research prototypes. Neural implicit and differentiable geometry methods are more mature in graphics, yet their integration into CAE workflows is just beginning. Nonetheless, the potential impact is significant. By improving geometry conditioning in non-CAD domains and replacing brittle heuristics in meshing algorithms, AI can reduce manual intervention, increase automation, and enable tighter coupling between meshing and simulation objectives. Future work is likely to explore AI-guided point insertion in tetrahedral meshing, prediction of refinement and coarsening based on error estimators, and differentiable formulations that embed meshing within simulation-driven optimization loops. Together, these advances point toward a future where unstructured mesh generation evolves from a heuristic bottleneck into a data-driven, adaptive process integrated seamlessly with defeaturing, classification, segmentation, and downstream FEA.
7 Block Structured Meshing
Many advanced physics solvers—such as those for shock physics or structural dynamics—depend on meshes composed entirely of quadrilaterals (2D) or hexahedra (3D). Solver accuracy and stability can be highly sensitive to mesh structure, making all-hex and block-structured approaches especially valuable [Benzley1995, Schneider2022]. The principle is straightforward: decompose a complex domain into four-sided or six-sided blocks and fill each with a structured grid, often using transfinite interpolation (TFI) [Thompson1998]. A widely adopted variant, the “pave-and-sweep” strategy [StatenCanannOwen1999] [RocaSarrate2010], generates a quadrilateral surface mesh and extrudes it through a third dimension to form hex elements.
Over decades, geometry- and topology-based methods such as mapping, submapping, and pave-and-sweep have provided the foundation of structured meshing. Production tools embody these strategies—Sandia’s Cubit® [Cubit15_2_UserDoc_2016] integrates an interactive environment for decomposition, while industrial packages like Gridgen [Chawner1995] and GridPro [Ferlemann2010] emphasize smoothing and topology-first multiblock generation for CFD. Yet despite their maturity, these methods leave the hardest step—discovering a valid block decomposition—largely in the hands of expert analysts. Progress in 2D has been tangible, but robust, general-purpose automation in 3D has remained elusive.
That humans can routinely identify good decompositions where algorithms fail points to what is missing: abstract reasoning, adaptive decision-making, and the ability to plan sequences of operations. Early attempts to inject AI into meshing, such as expert systems proposed by Lu, Gadh, and Tautges [lu2001feature], hinted at this direction but lacked the flexibility to generalize. Modern machine learning, and RL in particular, offers a more powerful framework. RL agents are explicitly designed to learn action sequences that optimize long-term goals, making them well aligned with the challenges of block decomposition, where the aim is to maximize completeness, mesh quality, and element conformity.
7.1 2D Block-Structured Meshing
While 2D block decomposition has direct applications, one of its roles today is as a proving ground for AI methods aimed at the more complex 3D case. DiPrete, Garimella, Cardona, and Ray [diprete2024reinforcement] recently reframed the problem using RL. Their agent learned to apply sequences of axis-aligned cuts to orthogonal polygons, decomposing them into quadrilateral blocks. The framework was built on Soft Actor–Critic (SAC) [haarnoja-2018-sac], whose maximum-entropy objective promotes exploration, stabilizes training, and supports continuous state and action spaces.
The state representation captured local geometry around vertices in fixed-length vectors, while the reward function encouraged quad generation with uniform geometric properties (e.g., area, aspect ratio) and lightly penalized unproductive actions. On axis-aligned datasets restricted to $X$- and $Y$-direction cuts, the agent consistently achieved $100\%$ quadrilateral decompositions. Follow-on work expanded the action set with bisections and generalized the approach to arbitrary polygons [RayRL2025]. For orthogonal shapes, the policy learned to favor edge-aligned cuts that produced rectangles; for general polygons, it generated many quadrilaterals but often left residual triangles. These results show that RL can capture effective decomposition strategies but that fully generalized behavior remains out of reach.
The limitations also highlight a key distinction from standard RL benchmarks: the meshing environment is not static but varies dramatically across CAD models. Capturing such variability with fixed-length features is brittle and often leads to noisy training. Promising directions include richer, multi-scale state representations, larger and more diverse training corpora, and the use of imitation learning [Argall2009SurveyLfD] to provide strong initial policies. In this way, 2D decomposition continues to serve not only as a useful capability in its own right but also as a critical stepping stone toward 3D block-structured and all-hex meshing, where the challenges—and the potential impact—are far greater.
An alternative to the approach taken in [diprete2024reinforcement] is to leave the geometry domain unchanged and build a block structure that discretizes it. This can be achieved by considering the work of Naranyan et. al. [Narayanan2024], who present a learning-based framework for improving the quality of unstructured triangular and quadrilateral meshes. Their approach learns to improve mesh quality via self-play RL with no prior heuristics. They use an actor-only version of the Proximal Policy Optimization (PPO) algorithm (see [schulman2017]) without a critic (i.e., value function) network. They demonstrate their agent’s ability to transform unstructured quadrilateral meshes—obtained by splitting triangles into four quads—into structured meshes that correspond to 2D block decomposition. Although the goal of [Narayanan2024] is not block decomposition, the applicability of their approach opens promising perspectives for pursuing this path.
Complementing these learning-based strategies, Paul et al. [PaulMCTS2025] frame 2D decomposition as Monte Carlo Tree Search (MCTS): starting from the CAD model as root, the search expands candidate partition operations while balancing exploitation and exploration. The method often finds near-optimal quadrilateral decompositions, showing that high-quality layouts can emerge without end-to-end training. Extending MCTS to 3D is especially promising—partial block layouts alone could guide experts and scaffold downstream learning. More broadly, search-based methods complement RL when reward shaping and state design are brittle, and they anticipate the demands of 3D all-hex and block-structured meshing: long-horizon reasoning under sweepability, feature-preservation, and quality constraints, developed next.
7.2 3D All-Hex and Block Structured
All-hex meshing of complex CAD models begins with a sound geometric decomposition. Downstream schemes, including unstructured approaches (for example, pave and sweep), structured mappings (for example, map), and block-structured layouts, then generate conforming, high-quality hexahedra. The central difficulty is not meshing a single sweepable or mappable block, but discovering a valid global decomposition and an executable sequence of operations that (i) respect sweepability constraints, (ii) preserve curvature and sharp features, and (iii) maintain element quality. Decades of practice have codified this workflow in tools such as Sandia’s Cubit®, which provide robust predicates and operations, yet the pivotal decomposition step remains largely manual. This gap motivates AI methods that can capture expert strategies, reason over long action horizons, and adapt decisions to complex geometry.
Recent work has begun to push in this direction. Patel et al. [Patel2021AutoHex] introduced Auto-Hex, which integrates geometric reasoning with RL in a closed loop with Cubit®. Starting from a boundary representation (B-Rep) and chordal-axis skeleton, the system identifies features such as T-junctions, sweep directions, and through-holes. An RL policy then proposes webcut actions conditioned on these cues, with feedback from the meshing environment guiding the agent to favor decompositions that yield hex or hex-dominant meshes. In related 2D work discussed earlier, DiPrete et al. [diprete2024reinforcement] used CAD shape testbeds to study RL-driven cut sequencing and reward design, while Zhang et al. [Zhang2025CADRLHex] frame 3D decomposition as a sequential decision problem and employ Soft Actor-Critic (SAC) to generate sweepable sub-volumes, showing promising decompositions on complex solids while highlighting open challenges in reward shaping and mesh quality.
In parallel, supervised learning has been explored as another route to automation. Quadros et al. [Quadros2025Bipartite] reduce a general 3D CAD model (outer B-Rep “skin”) to a 1D bipartite, heterogeneous graph derived from a 2D chordal-axis skeleton, then apply a graph neural network (GNN) to predict decomposition operations. Graph nodes carry geometric attributes (area, length, centroid), topological attributes (valency, interior loops), and meshing cues (principal sweep directions), with two-way mappings between the B-Rep, skeleton, and graph. Because labeled decompositions are costly, training is performed on synthetic CAD with action labels, yielding $98.72\%$ accuracy in predicting operations such as Cubit® webcuts. Current work is extending this approach to targeted CAD classes and coupling it with RL, as in Auto-Hex, to move beyond labeled datasets and learn the sequencing of webcut commands on general 3D geometries.
Taken together, these efforts show clear progress but also uneven maturity. In 2D, RL agents reliably decompose orthogonal polygons into quadrilaterals and show promising, if less consistent, performance on general shapes [diprete2024reinforcement]. In 3D, methods such as Auto-Hex [Patel2021AutoHex], SAC-based sequential decomposition [Zhang2025CADRLHex], and supervised GNN approaches [Quadros2025Bipartite] demonstrate feasibility on curated models and strong results on synthetic datasets, but remain at an early research stage and are not yet ready for routine use on complex industrial geometries.
The most promising path forward lies in hybridization and scale. Supervised models can provide strong priors for plausible decompositions, while RL agents can optimize their ordering and long-term consequences. Richer geometric encodings, larger and more diverse CAD corpora, and tighter coupling to production meshing environments will all be essential. Even partial automation—where algorithms propose cuts that analysts validate—could reduce manual effort from weeks to hours, broaden access to structured meshing, and improve solver robustness through higher-quality meshes. Looking further ahead, the integration of autonomous decomposition agents directly into meshing tools points toward the long-sought goal of true “push-button” all-hex meshing for classes of geometries that remain out of reach today.
8 Volumetric Parameterizations
Volumetric parameterization maps a solid 3D volume onto simple parametric domain (e.g., a cube or polycube). This mapping must be bijective and low-distortion to support applications such as hexahedral meshing, spline modeling and isogeometric analysis (IGA). While surface parameterization is well-established, extending these concepts to volumes introduces significant challenges due to higher dimensionality as well as complex geometry and topology.
Unlike block-structured meshing, which decomposes geometry into discrete blocks for local meshing, volumetric parameterization constructs a continuous global map from domain to volume. This enables two critical tasks: (i) generating high-quality all-hex meshes for simulation, and (ii) building tensor-product spline domains for IGA, where hex control meshes align naturally with spline-based discretizations [cottrell_isogeometric_2009].
Constructing smooth, bijective, and low-distortion volumetric maps remains an open problem. Complex geometries often preclude simple solutions, and existing methods struggle with element inversion, excessive distortion and extraordinary nodes. These issues directly impact mesh validity and the stability of numerical simulations, particularly for spline-based IGA.
Several frameworks target this problem from a meshing perspective. CubeCover [Nieser2011CubeCover] introduces volumetric frame fields and integer-grid maps to produce boundary-aligned parameterizations. Polycube methods deform the input into a polycube and apply regular grid pullbacks to extract hexahedral elements [Gregson2011Polycube, Fang2016ClosedFormPolycube]. Recent efforts improve robustness with intrinsic formulations [Mandad2022IntrinsicPolycubes] and quantization techniques [Brueckler2022Quantization], enhancing alignment and hex quality.
Beyond meshing, parameterization plays a central role in IGA and geometry processing. Early methods convert unstructured meshes into T-spline representations [wenyan2010a, Wenyan2011a]. Harmonic maps solve Laplace equations with Dirichlet boundaries to parameterize genus-0 and higher-genus volumes [zhang_solid_2012, Wenyan2013c, LeiLiu2012a]. Polycube decomposition has become standard for IGA [LZH2014, Liu2015, HZ2016CMAME, HZL2017], while advanced spline bases—e.g., weighted T-splines [LeiLiu2015a], truncated T-splines [ref:wei16a], and hierarchical B-splines [wei17a, wei2018blended]—are now supported in commercial solvers [Lai2016, Lai2017, yu2022hexgen]. Optimization-based methods introduce Jacobian regularization [Ji2022] and multi-patch schemes to ensure inter-domain continuity [Zou2025].
Despite these advances, robust parameterization for general CAD models remains elusive. Most pipelines still require expert input, difficulty in automation, suffer from high computational costs, and lack guarantees of bijectivity or low distortion—limiting mesh quality and analysis reliability.
Recent AI-driven methods offer a promising shift. Deep learning automates segmentation and feature alignment, transforming hours of manual work into seconds. DeepShape [DeepShapeCode] uses voxelized inputs and CNNs to learn mappings from geometry to polycube domains. DL-Polycube [yu2025dlpolycube] and DDPM-Polycube [yu2025DDPM] predict polycube structures using neural templates or diffusion models. Neural integral parameterization methods model the inverse map directly, maintaining bijectivity and handling boundary constraints [Zhan2025]. Reinforcement learning approaches [TONG2023102109] optimize parameterization strategies that preserve quality across complex inputs. Using genetic algorithms, Evocube [Dumery2022] focuses on the labelling problem, which aims to precompute polycube alignment by assigning one of the base axes to each boundary face of the input. When evaluated on a thousand smooth and CAD meshes, Evocube converges on accurate labels for a wide range of shapes.
Though these methods are still evolving, they demonstrate the feasibility of fully automated, and high-quality parameterization. Neural networks now produce bijective maps, generative models handle topological decomposition, and hybrid AI-classical methods combine learning with robust geometric frameworks.
The potential impact is transformative. Automation can drastically reduce modeling time, improve mesh quality, and lower the barrier for spline-based simulation at scale. While block-structured meshing partitions geometry, volumetric parameterization enables continuous mappings essential for high-fidelity simulation. The integration of AI with geometric theory represents the most scalable path forward.
9 Parallel Mesh Generation
As simulation fidelity and geometric complexity rise, large-scale FEA increasingly hinges on parallel mesh generation to sustain simulations with hundreds of millions of elements. Misconfigured runtimes can waste substantial wall time at this scale, so choosing the right parameters matters as much as the mesher itself. Both data and computation must be distributed over many processors in an HPC environment. Modern parallel meshers are indispensable, but they expose numerous runtime controls such as block size, work-partitioning strategy, NUMA affinity, and communication thresholds, and these choices have first-order effects on scalability and wall time. Selecting an optimal parameter vector $p$ is a nonlinear and architecture-sensitive problem that has traditionally depended on expert heuristics or costly trial and error.
Analytic performance models exist, but they typically rest on simplifying assumptions and generalize poorly across architectures and problem families [barker2005practical]. Empirical tuning is more faithful but costly: each evaluation consumes a full HPC submission, and queue delays often dominate turnaround time. Collecting even a few dozen labeled runs can take weeks in production settings [angelopoulos2018]. These constraints impede routine optimization when rapid results are needed.
Artificial intelligence offers a complementary path. In particular, supervised learning can train surrogate models that map parameter vectors to execution time [gamatie2019, ipek2005]. Training shifts cost up front. Once learned, the surrogate supports fast what-if exploration over unseen configurations without consuming additional HPC cycles, which enables wide synthetic searches of the parameter space. In what follows we focus on two complementary roles for AI: surrogates that predict performance and enable rapid configuration search, and learned controllers that replace fixed heuristics inside the parallel pipeline.
As a case study, [chrisochoides2025] trains compact feedforward surrogates for a general-purpose parallel mesher that predict execution time from a parameter vector $p$ using small real-world datasets. Two platforms were considered: 172 runs on a shared-memory cc-NUMA system [garner2025numa] and 68 runs on a multi-node distributed-memory system [garner2025dist]. A three-layer network per dataset was sufficient. The trained models then scored 100,000100{,}000 synthetic configurations in under $0.5$ s per case, turning weeks of batch exploration into minutes. On held-out tests the models reached MAPE $2.13\%$ on the shared-memory system and $5.68\%$ on the distributed-memory system [tsolakis2024, garner2025dist]. This accuracy was sufficient to select settings that reduced total execution time. Overall, surrogate-guided search cut tuning time-to-solution by more than $90\%$.
Beyond surrogates, recent work uses learning to decide how parallel meshing is organized. For load balancing in production workflows, trained classifiers select domain-decomposition settings and reduce time-to-solution compared to static heuristics, as shown for snappyHexMesh [Gangopadhyay2025ParCFD]. For decomposition, reinforcement learning with graph neural networks learns partitioners that compete with multilevel methods while exposing a trainable interface to work distribution and communication footprint [Gatti2022RLGNN]. For adaptation, adaptive mesh refinement has been posed as a sequential decision problem, and learned policies make locality-aware refinement choices that scale across partitions and generalize across PDEs and mesh sizes [Yang2023RLAMR]. Taken together, the results show that learned controllers can optimize decomposition, load balancing, and adaptive refinement, while surrogates accelerate search over the remaining parameters.
Positioned within this landscape, the surrogate-based study of [chrisochoides2025] addresses the parameter-tuning bottleneck by predicting execution time and enabling large synthetic searches that would otherwise consume weeks of HPC time. In parallel, other AI efforts intervene at structural decision points in the pipeline through learned load balancing [Gangopadhyay2025ParCFD], learned partitioning [Gatti2022RLGNN], and learning-based AMR [Yang2023RLAMR]. A near-term priority is portability for surrogate models beyond the training environment across geometry families and hardware by enriching inputs with quantitative geometric descriptors [foteinos2014, tsolakis2021] and by using transfer learning to adapt with minimal new data. A second priority is composition: pair surrogates with learned controllers to form an end-to-end AI-assisted workflow that partitions the domain, balances work, refines adaptively, and tunes runtime parameters under a single optimization loop. If realized, these advances would shift parallel meshing from a hand-tuned craft to a data-driven and reusable service embedded in large-scale simulation pipelines.
10 Scripting Automation
Preparing CAD models for simulation frequently relies on scripting as the standard practice. While some interactive operations can be performed ad hoc, most engineering simulations demand reproducibility and careful documentation. Scripts, often in the form of journal files containing Python or native CAD commands, serve this purpose by archiving the exact sequence of geometry operations applied to a model. This ensures that the process of simplifying and idealizing geometry can be repeated, audited, or adapted for future analyses.
Developing these scripts is rarely straightforward. Analysts typically discover effective command sequences through interactive trial and error: applying operations, attempting to mesh, revising geometry, and repeating until a successful workflow emerges. The difficulty arises because many geometry preparation problems require ordered sequences of operations rather than single isolated fixes. Defeaturing, for example, may involve multiple removals in a particular order. Block decomposition for hexahedral meshing requires long structured command sequences to partition geometry into meshable sub-blocks. Thin volume reduction [owen2024cad] demands careful transformations to collapse narrow solids into sheet bodies without breaking topology. Across all of these applications, the analyst’s final output is a script that embodies this iterative, experience-driven process.
Traditional approaches to scripting rely heavily on hand-crafted heuristics and expert intuition. Procedures are often encoded feature by feature, with analysts building libraries of macros or journal files that address narrow classes of problems. While effective in specific contexts, these methods are labor-intensive, brittle, and difficult to generalize across diverse geometries or simulation targets.
RL provides a natural way to address this challenge by framing script generation as a sequential decision-making problem. Unlike supervised methods that provide static recommendations, RL is designed to learn strategies that unfold over multiple steps. By treating CAD software as the environment, geometry operations as actions, and meshing or analysis outcomes as reward signals, RL agents can learn not only which operations to apply but also the order in which to apply them. Training on existing analyst scripts provides a foundation, while reward functions can encode expert priorities such as meshability, robustness, and fidelity to the target physics. In this way, RL has the potential to automate the generation of scripts that prepare CAD models for simulation, reducing manual effort and improving both efficiency and reproducibility.
Reinforcement learning methods for geometry preparation are implemented by casting CAD operations into the standard RL framework of environments, actions, and rewards. The environment is typically provided through an API to an existing geometry kernel, such as Cubit®  [CubitWebsite] or Open Cascade [OpenCascadeWebsite], which exposes the full set of modeling operations to the RL agent. Actions correspond to individual geometry commands that can be applied in sequence to modify the model, while the reward function encodes the analyst’s objectives, such as improving meshability or reducing unnecessary complexity. Training proceeds through trial-and-error interaction, allowing the agent to explore command sequences and gradually refine strategies that lead to high-quality model preparation scripts.
Recent work has demonstrated the viability of this approach in several application areas. In the context of block decomposition for hexahedral meshing, DiPrete et al. [diprete2024reinforcement] and Zhang et al. [zhang2025reinforcement] designed RL agents whose actions correspond to decomposition operations. The agents learn to select reference entities and cutting operations that subdivide complex geometries into meshable sub-blocks. The reward functions in these studies were crafted to encourage the formation of well-shaped sub-blocks with favorable surface and dihedral angles, while penalizing irregular cuts that degrade mesh quality. To support effective training, Zhang et al. [zhang2025reinforcement] also introduced a graph-based representation of CAD models. This representation maps B-rep entities to NURBS-based graph nodes, enabling the neural network to capture geometric and topological relationships. Using this setup, they reported hex-meshable volume ratios exceeding 95%, a significant milestone for automated decomposition.
Another example is the reduction of thin volumes in CAD models, a task that often requires ordered command sequences to collapse narrow solids into topologically valid sheet bodies. Owen et al. [owen2024cad] applied RL to this problem by defining actions that select thin volumes and determine appropriate sheet body placements. This approach preserved model topology while simplifying the geometry, and it is now being extended to automate defeaturing operations that improve tetrahedral mesh quality and reduce element counts. The long-term goal of this line of work is to provide analysts with automated script generation that mirrors expert decision-making, thereby reducing manual effort and accelerating the overall modeling-to-simulation workflow.
Although these early demonstrations show considerable promise, the maturity of RL methods for geometry preparation is still at an early stage. Most studies remain proofs of concept, focusing on constrained problem classes such as block decomposition or thin volume reduction. These results confirm that RL can, in principle, learn sequences of CAD operations that prepare models for simulation, but the methods have not yet been validated across the wide range of geometries encountered in real engineering practice. One notable exception is the thin volume reduction work of Owen et al. [owen2024cad], which has been integrated into Cubit® and is already providing useful capability. Even here, however, broader training data and wider analyst usage are needed before the approach can be considered mature.
The central challenge is that the outcome of interest is a script or procedure, but RL can only learn such procedures if the geometry is represented in a way that makes the effects of operations visible to the learning algorithm. In other words, the model must be able to detect differences in the geometry as it changes step by step, so that it can learn which sequences of operations lead to improved outcomes. Designing representations that expose these incremental changes is therefore critical.
Several strategies are being explored to address this issue. One approach is to convert B-rep entities into structured formats that can be processed by neural networks, such as sequence-based encodings [zhang2024brep2seq, zhang2024brepmfr]. Another is to use graph-based representations, where nodes and edges capture both geometry and topology, and graph neural networks (GNNs) are applied to process this information [wu2020comprehensive]. Graph-based methods are particularly promising because they allow the RL agent to reason about how operations modify geometry and topology in ways directly relevant to meshing and analysis. By combining these representation strategies with reinforcement learning, it may be possible to train agents that not only reproduce expert scripts but also discover new and more efficient procedures for geometry preparation.
11 Language-Driven Scripting
While reinforcement learning offers a framework for discovering sequences of CAD operations through trial and error interaction with geometry software, large language models (LLMs) [zhao2023survey] address a complementary aspect of the problem: the automation of CAD scripting through direct code synthesis. RL methods are designed to learn strategies over many iterations, discovering which procedures lead to successful outcomes. In contrast, LLMs are trained on large collections of text and code, enabling them to directly translate natural language prompts into executable scripts. Within the broader context of geometry preparation and meshing, these two approaches are not competing but rather mutually reinforcing. RL develops new procedural strategies, while LLMs provide an accessible interface for capturing, reusing, and deploying these strategies in a reproducible manner.
The motivation for LLM-driven scripting stems from the same challenges that have long motivated automation: CAD workflows require careful, reproducible documentation of geometry operations. Analysts traditionally create scripts manually, either through tedious trial and error in a graphical interface or by directly writing commands. While effective, this practice is slow, error-prone, and demands specialized expertise. LLMs reduce these barriers by allowing analysts to describe their goals in natural language. From these descriptions, the model generates CAD commands that can reproduce geometry, apply defeaturing, initiate meshing procedures, or prepare models for analysis. Importantly, the user remains in control: the generated code is transparent and can be inspected, refined, or overridden as needed.
Recent developments highlight several themes in how LLMs are being applied to CAD scripting. One line of work focuses on direct code generation. Fine-tuned models and database retrieval-augmentation frameworks have shown the ability to translate textual input into syntactically correct and semantically meaningful CAD code [R1, R4, R13]. These approaches often employ structured function calling or specialized annotation schemes that capture the semantics of CAD operations such as extrusions, fillets, and lofts [N1]. Iterative refinement loops, whether automated or assisted by user feedback, further improve robustness by detecting errors and resynthesizing code [R5, R6, R14].
A second direction expands beyond text to multimodal input. Systems now combine language with sketches, images, or point clouds, enabling workflows where a designer provides partial visual guidance that the model translates into complete parametric scripts. Examples include models trained on large datasets of CAD images paired with code [R13, R18, R26], as well as fine-tuned multimodal LLMs that directly edit boundary representations from text prompts while preserving geometric integrity [N2]. These methods move toward richer interactions where LLMs act as design assistants capable of interpreting intent expressed across multiple modalities.
A third theme emphasizes verification and correction. Because CAD scripting demands high fidelity, LLM-based workflows frequently incorporate feedback mechanisms. Some systems compare generated geometry against target specifications, iteratively correcting deviations [R5]. Others employ auxiliary vision–language models to check code execution and flag errors for resynthesis. These feedback loops echo the trial and error philosophy of RL, but they are carried out in a supervised or user-guided manner that accelerates convergence toward reliable scripts.
Together, these advances show that LLMs complement the procedural learning strengths of RL by providing a unifying interface that connects geometry representations, meshing technologies, and AI-driven methods discussed throughout this paper. Where RL agents may learn how to optimally defeature a model or decompose it for hexahedral meshing, an LLM can package such procedures into user-accessible scripts. Representations such as graph neural networks or B-rep encodings, originally motivated by the need to support RL training [zhang2024brep2seq, zhang2024brepmfr, wu2020comprehensive], can also support LLMs by enriching the context available for script generation.
Looking forward, the convergence of these technologies points toward the rise of AI agents [fang2025selfevolving] for CAD and meshing. Agents extend the capabilities of LLMs by combining language understanding with iterative reasoning, tool usage, and memory. In the context of geometry preparation, such agents could autonomously call geometry kernels, verify outputs, and refine operations, while engaging users through natural language. This unifies the strengths of both paradigms: RL provides the procedural foundation by discovering effective sequences of operations, LLMs make those procedures accessible by generating scripts from high-level prompts, and agent architectures orchestrate the entire process in a closed loop. The result is a vision of intelligent assistants that accelerate analysis, ensure reproducibility, and lower barriers to entry, while keeping analysts firmly in control of the workflow.
Recent advances indicate that LLM-driven CAD scripting systems perform reliably for routine design tasks [R9, R10, R13, R14]. Nonetheless, challenges persist. Models may struggle with highly constrained or complex geometries, and iterative refinement remains necessary to address error propagation and maintain high fidelity in the designs [R15, R17, R28, R30]. These limitations lead the literature to highlight areas that should be prioritizes as future research:
  • Expanding Domain-Specific and Multimodal Datasets: The lack of large-scale, high-quality CAD-specific and multimodal datasets (text, sketches, images, and CAD programs) limits the ability of LLMs to generalize across diverse design tasks. Future research should focus on curating and releasing broader datasets that capture complex geometries and real-world industrial contexts [R2, R9, R18].
  • Improving Spatial Reasoning and Geometric Fidelity: While LLMs can produce syntactically valid CAD scripts, they often fail on tasks requiring precise geometric or spatial reasoning. Developing solver-aided approaches, neurosymbolic frameworks, or enhanced spatial localization mechanisms could significantly improve fidelity for complex and constrained designs [R10, R23, R25].
  • Extending LLMs Beyond Geometry to Full Design-to-Manufacturing Pipelines: Research should go beyond geometry generation to encompass other stages of engineering workflows, including design evaluation, optimization, feature tagging for simulation/FEA, and integration with downstream manufacturing processes. This would enable seamless end-to-end AI-assisted design systems [R12, R19, R20].
In addressing these limitations, researchers can accelerate the evolution of LLMs to achieve a transformative shift in the digital design cycle. By automating the translation of natural language into precise and executable CAD code, these systems promise to accelerate the design process while lower the threshold for broad adoption.
12 Conclusion
This review set out with a simple premise: that artificial intelligence can serve as a facilitating technology for geometry preparation and meshing, complementing rather than replacing decades of foundational research in kernels, algorithms, and structured workflows. The survey has shown how that premise now manifests across the CAD-to-mesh pipeline. We examined methods for part classification and segmentation that automate the recognition of features within large assemblies, predictors that anticipate mesh quality before elements are generated, and AI-driven approaches to defeaturing that recommend cleanup operations. We reviewed early demonstrations of reinforcement learning for unstructured and block-structured meshing, new strategies for volumetric parameterizations, and surrogate modeling that accelerates parallel meshing. Finally, we highlighted how reinforcement learning and large language models can automate scripting, capturing analyst expertise in reproducible form.
Together, these advances show that AI is already reshaping how models are prepared for simulation, even if most methods remain at an early stage of maturity. This survey has highlighted representative technologies, illustrated how learning can extend established meshing practice, and pointed toward emerging opportunities across the pipeline. At the same time, it provides only a snapshot of a rapidly evolving field: important pieces are still missing, and the approaches described here will inevitably be refined, extended, and in some cases surpassed as new models, datasets, and workflows emerge.
Across these topics, several cross-cutting themes emerge. Most approaches face common bottlenecks: the lack of large curated datasets, standardized benchmarks, and robust geometry representations that can transfer across diverse CAD sources. The maturity of methods also varies widely. Classical machine learning for classification and mesh quality prediction has already entered production software such as Cubit®, delivering measurable efficiency gains. In contrast, reinforcement learning for meshing and volumetric parameterization remain at the prototype stage, while segmentation and defeaturing tools are intermediate—promising in research settings but not yet broadly deployed. The most promising direction is integration: connecting classification, prediction, and defeaturing into closed-loop workflows; combining supervised and reinforcement learning; and leveraging large language models for scripting and natural interfaces. Progress in the next five years will depend less on isolated algorithmic advances than on building shared datasets, evaluation standards, and hybrid pipelines that unify AI with established meshing practice.
In this spirit, the review is not intended as a final word but as a springboard. We hope it serves as both a resource for practitioners seeking to incorporate AI into their workflows and a guidepost for researchers advancing the next generation of tools. The open challenges ahead, including richer data curation, standardized evaluation, B-rep-native learning, uncertainty-aware decision-making, and orchestration across the full simulation pipeline, offer fertile ground for continued exploration. By capturing current progress while pointing toward what lies ahead, we aim to support innovation across the community and invite future contributions that will ultimately advance and surpass the capabilities described here, bringing us closer to overcoming the long-standing CAD-to-mesh bottleneck in simulation.
[…]
Acknowledgments Sandia National Laboratories is a multimission laboratory managed and operated by National Technology & Engineering Solutions of Sandia, LLC, a wholly owned subsidiary of Honeywell International Inc., for the U.S. Department of Energy’s National Nuclear Security Administration under contract DE-NA0003525.